Please help me solve this problem.

Yu-Feng Liang bought a used car for $6500. He made a down payment of $1000 and paid off the balance in 48 monthly payments of $159 each.
What was the finance charge?

Answer by mananth(16946)

Yu-Feng Liang bought a used car for $6500. He made a down payment of $1000 and paid off the balance in 48 monthly payments of $159 each.
..
$6500-$1000 = $5,500
48 months $159
Total payment = 159*48= $7632
..
Borrowing was $6500
difference = interest
7632-5500= 2132
.
r= interest/(p*t)
r=2132/(5500*4)
r= 0.0969
= 9.7 percent simple interest
...
m.ananth@hotmail.ca